Find the root of the equation tanx + tanhx = 0 which lies in the interval (1.6, 3.0) correct to four significant digits using the method of false position.
1
Expert's answer
2020-10-22T17:25:15-0400

Given "f(x)=\tan x+\tanh x"

"f(a)=f(1.6)=\tan(1.6)+\tanh(1.6)=-33.3109"

"f(b)=f(3)=\tan(3)+\tanh(3)=0.8525"

Using iterative Regula-Falsi Formula, the first approximation,


"x_1=a-\dfrac{b-a}{f(b)-f(a)}f(a)=""=1.6-\dfrac{3-1.6}{0.8525+33.3109}(-33.3109)=2.9324"



"f(2.9324)=0.7821"


"a=1.6, b=2.9324"


"x_2=1.6-\dfrac{2.9324-1.6}{0.7821+33.3109}(-33.3109)=2.9018"

"f(2.9018)=0.7495"


"a=1.6, b=2.9018"


"x_3=1.6-\dfrac{2.9018-1.6}{0.7495+33.3109}(-33.3109)=2.8732"

"f(2.8732)=0.7186"


"a=1.6, b=2.8732"


"x_4=1.6-\dfrac{2.8732-1.6}{0.7186+33.3109}(-33.3109)=2.8463"

"f(2.8463)=0.6891"


"a=1.6, b=2.8463"


"x_5=1.6-\dfrac{2.8463-1.6}{0.6891+33.3109}(-33.3109)=2.8211"

"f(2.8211)=0.6609"

"..."


"x_{149}=2.36550"

"x_{150}=2.36548"

"x_{151}=2.36546"

"..."


"x_{499}=2.36502"

Root of "f(x)" is "2.365"